Approximately 3% of the newboms are twins. Chocse 5 newborns at random. Mhat is
the probability that none twins? What is the probability that at least 1 is a twin
?
Answer #1

(a)

Question: What is the probability that none twins?

Answer:
Binomial Distribution

n = 5

p = 0.03

q = 1 -p = 0.97

P(X=0) = qn = 0.975 = 0.8587

So, answer is:

0.8587

(b)

Question: What is the probability that at least 1 is a twin

Answer:
P(X>0) = 1 - P(X=0)

        = 1 - 0.8587

      = 0.1413

So,

Answer is:

0.1413